1. Will a 0.30 m × 0.30 m × 0.30 m empty copper (ρCu = 8890 kg/m3) box float in water? The total mass of copper is 3.2 kg.
1
Expert's answer
2022-04-18T15:18:46-0400

The average density of a box

"\rho=\frac{m}{V}=\frac{3.2\rm \: kg}{(0.3\:\rm m)^3}=119\:\rm kg/m^3"

Since

"\rho<\rho_w=1000\;\rm kg/m^3"

hence the box will float in water.